# Verifiable mission planning for space operations

### v3

@@ Abstract
Spacecraft must operate under environmental and actuator uncertainties while meeting strict safety requirements. Traditional approaches rely on scenario-based heuristics that fail to account for stochastic influences, leading to suboptimal or unsafe plans. We propose a finite-horizon, chance-constrained Markov decision process for mission planning, where states represent mission and vehicle parameters, actions correspond to operational adjustments, and temporal logic specifications encode operational constraints. We synthesize policies that optimize mission objectives while ensuring constraints are met with high probability. Applied to the GRACE-FO mission, the approach accounts for stochastic solar activity and uncertain thrust performance, yielding maneuver schedules that maximize scientific return and provably satisfy safety requirements. This work demonstrates how Markov decision processes can be applied to space missions, enabling autonomous operation with formal guarantees.
@@ 1 Introduction
Planning spacecraft operations is critical for sustaining mission objectives, as missions must adapt to variable conditions and limited resources. Traditional approaches rely on precomputed scenarios or heuristic scheduling rules that assume fixed environmental conditions[1]. These methods often fail to capture the stochastic nature of orbital decay and space weather variability, leading to inefficient fuel use or violations of operational constraints.
@@
The GRACE-FO mission provides a motivating case study. Designed to map Earth’s gravity field through tandem satellite operations, it has faced two major challenges: malfunctioning thrusters limiting orbit control, and the failure of one satellite’s accelerometer. To mitigate the resulting degradation in measurement accuracy, the mission must operate at higher altitudes, where atmospheric drag is reduced. This requirement, combined with uncertain solar activity and fuel limitations, makes orbit planning particularly difficult. Previous orbit management strategies, such as those of Save et al. [1], rely on fixed solar cycle scenarios; however, they do not capture variability in solar flux and can result in premature reentry under adverse conditions.
@@
We address these challenges by formulating spacecraft mission planning as a finite-horizon Markov decision process (MDP), a framework for sequential decision making under uncertainty [2]. For GRACE-FO, states capture mission-relevant quantities such as orbital altitude and fuel reserves, actions represent discrete maneuver options, and transition probabilities model the effects of uncertain factors like solar flux and thrust variability. The reward function encodes mission objectives, such as maintaining observation quality at higher altitudes and avoiding repeated groundtrack orbits. Solving the MDP yields maneuver policies that maximize mission return while enforcing constraints such as minimum altitude and maneuver spacing. Unlike traditional scenario-based methods, this approach directly incorporates uncertainty into the planning model, enabling adaptive maneuver scheduling with formal, probabilistic safety guarantees.
Prior studies on mission planning under uncertainty have approached the problem from several directions. Scenario-based analyses, such as those of Save et al. [1], evaluate a fixed set of worst-case, nominal, and best-case conditions, but this often produces overly conservative strategies. Decision-theoretic approaches have sought to improve flexibility: Eddy et al. [3] applied MDPs to integrate uncertainty into satellite task planning, enabling trade-offs across multiple objectives. Other work has pursued optimization-based methods, such as Isaji and Ho [4], who combined stochastic programming with decomposition techniques to jointly address mission planning and spacecraft design. In parallel, robust design strategies have been explored, including evidence-theory formulations for quantifying uncertain parameters [5] and risk-averse planning techniques that incorporate coherent risk measures to mitigate adverse outcomes [6]. While these studies advance uncertainty-aware mission planning, they do not provide explicit probabilistic safety guarantees tailored to spacecraft operations, the focus of this work.
We develop a finite-horizon MDP method for probabilistic mission planning and safety verification in spacecraft operations. The model embeds environmental and system uncertainties directly into state transitions, enabling maneuver strategies that respect formal operational constraints. We apply this approach to the GRACE-FO mission, demonstrating how the resulting policies improve fuel efficiency and data quality while satisfying safety requirements such as minimum altitude. By extending scenario-based orbit management [1] and advancing decision-theoretic planning [3], our approach presents a proof of concept for applying formal safety guarantees to space mission planning, pointing toward certifiable autonomy in future missions.
@@ 2 Probabilistic Policy Synthesis and Safety Verification Using Markov Decision Processes
We model spacecraft decision-making under uncertainty as a finite-horizon Markov decision process (MDP). An MDP is a common paradigm for modeling planning and acting in stochastic environments with nondeterministic action selection [2]. For mission planning, we focus on finite-horizon MDPs, wherein the decision-making process is performed over a finite number of time steps called a planning horizon. In the context of a spacecraft mission, the environment stochasticity could be due to, e.g., uncertainties in future space weather conditions, while actions could correspond to performing altitude raises. Similarly, the planning horizon could be the operational lifetime of the mission, while individual time steps are specific points in time at which actions are chosen, e.g., daily, weekly, or monthly.
@@
To model the safety specification $\varphi$, we can characterize it in terms of a temporal logic formula [7, 8]. The foundational building blocks of a temporal logic formula are atomic propositions, which evaluate to either true or false, e.g., “the spacecraft has nonzero fuel reserves remaining.”
Note that it is straightforward to generalize the constraint (13b) to the probability of multiple mission specifications. We refer the reader to Forejt et al.[9] section 8 for further details.
@@ 3.1 Orbital Mechanics
in which $\dot{\theta}$ is the rotation rate of the Earth. The atmospheric density $\rho_{A}$ is evaluated using the NRLMSISE-00 model [10, 11, 12] as a function of altitude, and the solar flux at 10.7 cm (F10.7) obtained from NOAA predictions [13].
In tandem with the preceding Cartesian representation of the spacecraft dynamics, one can instead characterize its orbit in terms of the Keplerian orbital elements $\{a,e,i,\omega,\Omega,f\}$, where $a$ is the semimajor axis length, $e\in[0,1)$ is the eccentricity, $i$ is the inclination, $\omega$ is the argument of periapse, $\Omega$ is the longitude of the ascending node, and $f$ is the true anomaly [14]. These orbital elements are visualized in Figure 2, where the angles are defined with respect to the Earth-centered inertial reference frame.
@@
We refer the reader to [14] for information on how to convert between the Cartesian and Keplerian representations of the spacecraft state.
@@ 3.2 Space Weather
Due to the atmospheric drag force $F_{\mathrm{drag}}$, the eccentricity and semimajor axis of the spacecraft orbit decrease monotonically over its operational lifetime, eventually approaching values of $R_{e}$ (Earth’s radius) and zero, respectively. As a result, the spacecraft gradually descends to a circular orbit of zero altitude [14]. The extent and variability of this degradation depend heavily on the atmospheric density at the spacecraft’s altitude, which is strongly influenced by solar activity.
The solar cycle governs the Sun’s activity levels and, in turn, the atmospheric environment encountered by the spacecraft. However, the duration and intensity of individual cycles vary; larger cycles tend to last longer than 11 years, while smaller cycles are typically shorter. Solar activity is commonly categorized into three levels: high, moderate, and low [1]. During high solar activity, increased ultraviolet and X-ray radiation heat and expand the Earth’s upper atmosphere, raising atmospheric density at orbital altitudes and increasing drag forces. This accelerates orbital decay, requiring more frequent station-keeping maneuvers and increasing fuel consumption. In moderate activity periods, atmospheric density variations are less extreme, allowing for more stable orbit maintenance. Conversely, low solar activity leads to reduced drag, minimizing fuel expenditure and extending mission lifetime.
@@ 4.1 GRACE-FO mission
The Gravity Recovery and Climate Experiment (GRACE), launched in 2002, was a NASA mission that sought to obtain high-accuracy estimates of Earth’s gravity field from twin satellites in near-polar, near-circular orbits [15]. The GRACE-FO (Gravity Recovery and Climate Experiment Follow-On) mission, launched in May 2018, continues the legacy of the original GRACE mission to measure Earth’s gravity field with unprecedented precision. By monitoring variations in the inter-satellite range using K-Band Ranging and Laser Ranging Interferometer technologies, GRACE-FO provides critical data on global surface mass changes.
One major challenge arises from the failure of the accelerometer on GRACE-FO 2 (GF2), rendering it unable to measure non-gravitational forces directly. To compensate, the data from the accelerometer of GRACE-FO 1 (GF1) are transplanted to GF2, relying on the assumption that both satellites experience nearly identical environmental conditions within the $\sim 24–30$ s separation delay. The transplant method recovers differential non-gravitational forces with high fidelity[16] under nominal conditions, but it introduces signal-proportional errors that can reach a few percent of the drag signal in high-drag regimes, which degrades the resulting gravity fields.
@@ 4.3 Reward Modeling
At a high level, the objective of the GRACE-FO mission is to provide high-quality observations of changes in Earth’s gravitational field [17]. The quality of these observations generally increases with lower atmospheric density, which occurs at higher altitudes. Due to this fact, the reward function $R$ should be independent of the choice of the action selection and only a function of the current state (i.e., altitude) and time step of the mission. The dependence of the reward function $R$ on time could be due to, e.g., time-varying atmospheric conditions that affect the observation quality. Thus, we introduce $R(s,:,h)\triangleq R(s,a,h)$ for all $a\in\mathcal{A}$ to denote the reward (a function of the observation quality) for being in a state $s\in\mathcal{S}$ corresponding to a fixed, discrete altitude. In other words, $R(s,:,h)$ is independent of the action taken. Note that $R(s,:,h)$ should be non-negative and monotonically increasing as the corresponding altitude of $s$ increases.
An additional hindrance to observation quality is the existence of resonant orbits, which are orbits with a periodic ground track [18]. Due to the periodicity of this ground track, less of Earth’s surface is observed, reducing the quality of the global gravitational field estimation. Thus, for any discrete altitude state $s\in\mathcal{S}_{a}$ for which a resonant orbit exists in its corresponding altitude band $s-\frac{R_{a}}{2},s+\frac{R_{a}}{2}]$, we can set $R(s,:,h)=0$. Alternatively, given the periodicity of the resonant orbit $i$, $res_{i}$ (i.e., the rate at which the ground track repeats), one can scale the reward function by a multiplicative factor $\alpha<1$ to account for variations in the reduction of observation quality. The reward function can then be defined as
@@
Since the mean motion $n$ is directly related to the semi-major axis via Kepler’s law, this equation is solved iteratively using the Lagrange Planetary Equations with Klokočník’s method [19]. We then obtain the resonant altitudes. The resulting normalized repetition rate is then incorporated into the reward function (11).
@@ 5.1 Deterministic Planning
In the first set of simulations, the solar flux is assumed to be fixed at its expected value (as provided by NOAA[13]), and the cost for performing an orbit-raising maneuver is treated as deterministic. With these assumptions, the MDP is solved using a backward-recursion scheme. Figure 7 illustrates the deterministic evolution of altitude and fuel consumption. In this simulation, the spacecraft’s altitude increases in steps, reaching a final altitude slightly below 450 km, while maneuvers occur approximately every 6 months. These results indicate that the solution consistently maintains the spacecraft above 300 km.

### v4

@@ Abstract
Spacecraft must operate under environmental and actuator uncertainties while meeting strict safety requirements. Traditional approaches rely on scenario-based heuristics that fail to account for stochastic influences, leading to suboptimal or unsafe plans. We propose a finite-horizon, chance-constrained Markov decision process for mission planning, where states represent mission and vehicle parameters, actions correspond to operational adjustments, and temporal logic specifications encode operational reach–avoid constraints. We synthesize policies that optimize mission objectives while ensuring constraints are met with high probability. Applied to the GRACE-FO mission, the approach accounts for stochastic solar activity and uncertain thrust performance, yielding maneuver schedules that maximize scientific return and provably satisfy safety requirements. We demonstrate how Markov decision processes can be applied to space missions, enabling autonomous operation with formal guarantees.
@@ 1 Introduction
Planning spacecraft operations is critical for sustaining mission objectives, as missions must adapt to variable conditions and limited resources. Current in-operation approaches rely on precomputed scenarios or heuristic scheduling rules that assume fixed environmental conditions[save2023gracefo]. These methods often fail to capture the stochastic nature of orbital decay and space weather variability, leading to inefficient fuel use or violations of operational constraints.
@@
The GRACE-FO mission provides a motivating case study. Designed to map Earth’s gravity field through tandem satellite operations, it has faced two major challenges: malfunctioning thrusters limiting orbit control, and the failure of one satellite’s accelerometer. To mitigate the resulting degradation in measurement accuracy, the mission must operate at higher altitudes, where atmospheric drag is reduced. This requirement, combined with uncertain solar activity and fuel limitations, makes orbit planning particularly difficult. Previous orbit management strategies for GRACE-FO, such as Save et al. [save2023gracefo] work, rely on fixed solar cycle scenarios and do not capture variability in solar flux, which can result in premature reentry under adverse conditions.
@@
We address these challenges by formulating spacecraft mission planning as a finite-horizon Markov decision process (MDP), a framework for sequential decision making under uncertainty [puterman2014markov]. For GRACE-FO, states capture mission-relevant quantities such as orbital altitude and fuel reserves, actions represent discrete maneuver options, and transition probabilities model the effects of uncertain factors like solar flux and thrust variability. The reward function encodes mission objectives, such as maintaining observation quality at higher altitudes and avoiding repeated groundtrack orbits. Solving the MDP yields maneuver policies that maximize mission return while enforcing constraints such as minimum altitude and maneuver spacing. Unlike traditional scenario-based methods, this approach directly incorporates uncertainty into the planning model, enabling adaptive maneuver scheduling with formal, probabilistic safety guarantees.
Prior studies on mission planning under uncertainty have approached the problem from several directions. Scenario-based analyses, such as the study by Save et al. [save2023gracefo], evaluate a fixed set of worst-case, nominal, and best-case conditions, but this often produces overly conservative strategies that lack responsiveness to evolving uncertainties and cannot adapt when conditions deviate from those cases. Another approach is to use stochastic optimization and control techniques [chai2019stochastic, echigo2025autonomy, ozaki2020tube]. These methods typically operate over short horizons and emphasize local constraint handling, making them less suitable for multi-year, high-level mission planning with operational constraints. At longer horizons, decision-rule synthesis has been used to enable flexible logistics policies under uncertainty [chen2021flexibility]. However, this approach targets resource allocation rather than closed-loop planning and does not enforce formal safety constraints. In contrast, MDPs provide a sequential decision-making model that naturally supports uncertainty, policy synthesis, and enforcement of discrete safety or timing rules. Prior applications of MDPs include satellite task planning, maneuver scheduling, and resource management under uncertainty [eddy2020markov, kuhl2025markov, rommel2025optimal, hibbard2025autonomous]. However, these efforts generally encode constraints heuristically or via soft penalties, optimizing expected performance without certifying that safety-critical constraints are satisfied with high probability. Temporal logic has been used for formal specification [hibbard2023trajectory], but automaton-based methods require forming a product between the MDP state space and the specification automaton (and, for time-varying rules, an explicit time index), which increases complexity and limits scalability for long-term planning.
We propose a finite-horizon, time-dependent MDP for probabilistic mission planning in spacecraft operations, embedding environmental and system uncertainties. Operational rules are encoded as reach–avoid constraints, and safety is verified via backward dynamic programming on the state–time space, computing satisfaction probabilities without a specification-automaton product [hibbard2023trajectory, baier2008principles, forejt2011automated]. We apply this approach to GRACE-FO, demonstrating policies that meet mission objectives while satisfying operational safety requirements (e.g., minimum altitude). This extends scenario-based orbit management [save2023gracefo] and advances decision-theoretic planning for spacecraft operations [eddy2020markov], providing formal safety guarantees in space mission planning.
@@ 2 Probabilistic Policy Synthesis and Safety Verification Using Markov Decision Processes
We model spacecraft decision-making under uncertainty as a finite-horizon Markov decision process (MDP). An MDP is a common paradigm for modeling planning and acting in stochastic environments with nondeterministic action selection [puterman2014markov]. For mission planning, we focus on finite-horizon MDPs, wherein the decision-making process is performed over a finite number of time steps called a planning horizon. In the context of a spacecraft mission, the environment stochasticity could be due to, e.g., uncertainties in future space weather conditions, while actions could correspond to performing altitude raises. Similarly, the planning horizon could be the operational lifetime of the mission, while individual time steps are specific points in time at which actions are chosen, e.g., daily, weekly, or monthly.
@@
To model the safety specification $\varphi$, we can characterize it in terms of a temporal logic formula [baier2008principles, rozier2011linear]. The foundational building blocks of a temporal logic formula are atomic propositions, which evaluate to either true or false, e.g., “the spacecraft has nonzero fuel reserves remaining.”
Note that it is straightforward to generalize the constraint (13b) to the probability of multiple mission specifications. We refer the reader to Forejt et al.[forejt2011automated] section 8 for further details.
@@ 3.1 Orbital Mechanics
in which $\dot{\theta}$ is the rotation rate of the Earth. The atmospheric density $\rho_{A}$ is evaluated using the NRLMSISE-00 model [picone2002nrlmsise, emmert_nrlmsis_2021, emmert_nrlmsis_2022] as a function of altitude, and the solar flux at 10.7 cm (F10.7) obtained from NOAA predictions [miesch2024solar].
In tandem with the preceding Cartesian representation of the spacecraft dynamics, one can instead characterize its orbit in terms of the Keplerian orbital elements $\{a,e,i,\omega,\Omega,f\}$, where $a$ is the semimajor axis length, $e\in[0,1)$ is the eccentricity, $i$ is the inclination, $\omega$ is the argument of periapse, $\Omega$ is the longitude of the ascending node, and $f$ is the true anomaly [prussing1993orbital]. These orbital elements are visualized in Figure 2, where the angles are defined with respect to the Earth-centered inertial reference frame.
@@
We refer the reader to [prussing1993orbital] for information on how to convert between the Cartesian and Keplerian representations of the spacecraft state.
@@ 3.2 Space Weather
Due to the atmospheric drag force $F_{\mathrm{drag}}$, the eccentricity and semimajor axis of the spacecraft orbit decrease monotonically over its operational lifetime, eventually approaching values of $R_{e}$ (Earth’s radius) and zero, respectively. As a result, the spacecraft gradually descends to a circular orbit of zero altitude [prussing1993orbital]. The extent and variability of this degradation depend heavily on the atmospheric density at the spacecraft’s altitude, which is strongly influenced by solar activity.
The solar cycle governs the Sun’s activity levels and, in turn, the atmospheric environment encountered by the spacecraft. However, the duration and intensity of individual cycles vary; larger cycles tend to last longer than 11 years, while smaller cycles are typically shorter. Solar activity is commonly categorized into three levels: high, moderate, and low [save2023gracefo]. During high solar activity, increased ultraviolet and X-ray radiation heat and expand the Earth’s upper atmosphere, raising atmospheric density at orbital altitudes and increasing drag forces. This accelerates orbital decay, requiring more frequent station-keeping maneuvers and increasing fuel consumption. In moderate activity periods, atmospheric density variations are less extreme, allowing for more stable orbit maintenance. Conversely, low solar activity leads to reduced drag, minimizing fuel expenditure and extending mission lifetime.
@@ 4.1 GRACE-FO mission
The Gravity Recovery and Climate Experiment (GRACE), launched in 2002, was a NASA mission that sought to obtain high-accuracy estimates of Earth’s gravity field from twin satellites in near-polar, near-circular orbits [tapley2004gravity]. The GRACE-FO (Gravity Recovery and Climate Experiment Follow-On) mission, launched in May 2018, continues the legacy of the original GRACE mission to measure Earth’s gravity field with unprecedented precision. By monitoring variations in the inter-satellite range using K-Band Ranging and Laser Ranging Interferometer technologies, GRACE-FO provides critical data on global surface mass changes.
One major challenge arises from the failure of the accelerometer on GRACE-FO 2 (GF2), rendering it unable to measure non-gravitational forces directly. To compensate, the data from the accelerometer of GRACE-FO 1 (GF1) are transplanted to GF2, relying on the assumption that both satellites experience nearly identical environmental conditions within the $\sim 24–30$ s separation delay. The transplant method recovers differential non-gravitational forces with high fidelity[harvey2024graced] under nominal conditions, but it introduces signal-proportional errors that can reach a few percent of the drag signal in high-drag regimes, which degrades the resulting gravity fields.
@@ 4.3 Reward Modeling
At a high level, the objective of the GRACE-FO mission is to provide high-quality observations of changes in Earth’s gravitational field [landerer2020extending]. The quality of these observations generally increases with lower atmospheric density, which occurs at higher altitudes. Due to this fact, the reward function $R$ should be independent of the choice of the action selection and only a function of the current state (i.e., altitude) and time step of the mission. The dependence of the reward function $R$ on time could be due to, e.g., time-varying atmospheric conditions that affect the observation quality. Thus, we introduce $R(s,:,h)\triangleq R(s,a,h)$ for all $a\in\mathcal{A}$ to denote the reward (a function of the observation quality) for being in a state $s\in\mathcal{S}$ corresponding to a fixed, discrete altitude. In other words, $R(s,:,h)$ is independent of the action taken. Note that $R(s,:,h)$ should be non-negative and monotonically increasing as the corresponding altitude of $s$ increases.
An additional hindrance to observation quality is the existence of resonant orbits, which are orbits with a periodic ground track [pie2008subcycle]. Due to the periodicity of this ground track, less of Earth’s surface is observed, reducing the quality of the global gravitational field estimation. Thus, for any discrete altitude state $s\in\mathcal{S}_{a}$ for which a resonant orbit exists in its corresponding altitude band $s-\frac{R_{a}}{2},s+\frac{R_{a}}{2}]$, we can set $R(s,:,h)=0$. Alternatively, given the periodicity of the resonant orbit $i$, $res_{i}$ (i.e., the rate at which the ground track repeats), one can scale the reward function by a multiplicative factor $\alpha<1$ to account for variations in the reduction of observation quality. The reward function can then be defined as
@@
Since the mean motion $n$ is directly related to the semi-major axis via Kepler’s law, this equation is solved iteratively using the Lagrange Planetary Equations with Klokočník’s method [klokovcnik2003fine]. We then obtain the resonant altitudes. The resulting normalized repetition rate is then incorporated into the reward function (11).
@@ 5.1 Deterministic Planning
In the first set of simulations, the solar flux is assumed to be fixed at its expected value (as provided by NOAA[miesch2024solar]), and the cost for performing an orbit-raising maneuver is treated as deterministic. With these assumptions, the MDP is solved using a backward-recursion scheme. Figure 7 illustrates the deterministic evolution of altitude and fuel consumption. In this simulation, the spacecraft’s altitude increases in steps, reaching a final altitude slightly below 450 km, while maneuvers occur approximately every 6 months. These results indicate that the solution consistently maintains the spacecraft above 300 km.